### v1

@@ 5.1 Visual Question Answering
Hence we evaluate the effectiveness of FlexCap-LLM on several image VQA benchmarks such as OKVQA [33], GQA [21], and VizWiz [16], and video question answering benchmarks such as MSRVTT [51] and MSVD [50]. Diverse characterictics of these datasets helps gaining better insight on FlexCap’s capabilities. We report the commonly used accuracy metric for each dataset.
@@
(a) GQA results on test-dev set
@@
FlexCap-LLM 37.1
(c) VizWiz results on val set
@@
Method MSRVTT-QA MSVD-QA
@@
Table 2: Zero-shot video question answering results reported on MSRVTT-QA and MSVD-QA on test set. FlexCap-LLM is better than other zero-shot baselines for video VQA benchmarks.
5.1.1 Image Question Answering.
First we evaluate FlexCap-LLM on GQA, OKVQA and VizWiz image VQA benchmarks in a zero-shot setting, meaning that our approach is not trained with the task or the corresponding dataset. The results on these benchmarks are presented in Table 1.
Compositional VQA. The GQA dataset is for evaluating the performance on complex compositional questions. As FlexCap produces information for multiple visual elements in the scene with their corresponding locations, it is quite well-suited for questions on compositional understanding of the image. On this benchmark, as shown in Table 0(a), FlexCap-LLM outperforms all the recent baselines.
VQA with External Knowledge. OKVQA dataset is particularly designed for evaluating the ability to answer questions about images that require external knowledge which is not readily available on the image. Hence it requires multiple levels of understanding of the content, and reasoning with that information, which is well-suited for applying FlexCap. In Table 0(b) we show our performance on OKVQA is superior to strong baselines such as Flamingo and ViperGPT which highlights the effectiveness of the mix of generic and specific descriptions generated by FlexCap. Unlike other baselines which use the question, FlexCap generates the captions without having access to the question.
VQA with atypical images. We also evaluate on VizWiz, which contains visual questions asked by people who are visually impaired. Unlike web content, in these images the objects and the scene are not always well-centered, hence this dataset contains many out-of-distribution samples compared to typical web-crawled datasets. We report the results of this experiment in Table 0(c). Nevertheless, our approach significantly outperforms Flamingo [2] which also reports zero-shot performance on this dataset.
5.1.2 Video Question Answering.
We also evaluate FlexCap-LLM on zero-shot video question answering datasets MSRVTT-QA and MSVD-QA [50]. The results on these benchmarks are presented in Table 2. For processing the video, we sample 8 frames uniformly from the video. We pass each of these frames through FlexCap to produce captions of objects and regions. We then combine all the object captions from the different frames into one prompt for the LLM.
We observe FlexCap-LLM significantly exceeds the performance of the Flamingo model in the zero-shot setting. These results highlight the zero-shot effectiveness of our method, which can solve tasks in the video domain even though both FlexCap and the LLM were not trained for those tasks.
5.2 Dense Captioning
Fine-tuning FlexCap. We fine-tune the pre-trained FlexCap model on the Visual Genome train split for 60k steps with a lower learning rate of $61e-$ 6 at a resolution of $448\times 448$.
Captioning GT boxes. Following the evaluation procedure from [25], we evaluate captioning of the ground-truth boxes in Visual Genome. Since this setting removes the localization task, we have a cleaner evaluation of only the region captioning problem. The results of this experiment are provided in Table 3 in which we show that FlexCap achieves better performance compared to other approaches evaluated in this setting.
Captioning GRIT boxes. In this experiment, we want to compare against other approaches that perform both localization and captioning. We are measuring how well FlexCap performs when provided with box proposals. Table 3 shows FlexCap obtains better performance compared to other approaches evaluated in this setting. Since, in our work we do not propose any localization module we use GRIT’s [49] region proposals as the input boxes for our model. This also allows us to directly compare our captioning capabilities against GRIT. We find that our approach outperforms GRIT at this task even though we test at a lower resolution of $448\times 448$.
@@ 5.3 Open-Ended Object Detection
Open-ended object detection, that is identifying all the objects in the image with rich descriptions, can be achieved in two ways: (a) localize-then-describe or (b) describe-then-localize. With FlexCap, we employ a localize-then-describe approach where boxes are generated with a box proposal mechanism (i.e. OWL-ViT) and then described by FlexCap. We compare these results with a describe-then-localize approach where we first ask a state-of-the-art (SOTA) VLM (i.e. LLAVA [32]) for a comprehensive description of the image with all objects and then localize these descriptions using a SOTA open-vocabulary object detection method (i.e. OWL-ViT [36]). The Recall of all the objects in the image is a major criteria in open-ended object detection, and as shown in Fig. 4(e)(a), we demonstrate that localize-then-describe approach powered with FlexCap performs significantly better compared to describe-then-localize approach using existing works.
In Fig. 4(e) (b-e), we observe that our localize-then-describe approach retrieves more parts of the image, particularly the small and medium size objects, compared to the describe-then-localize approach. This is mainly because most VLMs are trained to describe the salient objects in the image rather than exhaustively list all the objects present in the scene. We show some typical examples of missed detections in Fig. 6.
@@
Figure 9: Diverse captioning with Prefixes. FlexCap can be used to perform conditional captioning of images. In all three examples, the input bounding box covers the whole image. We show three prefixes: "a", "the photo was taken in ", "notice" resulting in diverse captions for the same image.
5.4 Qualitative Results
@@
Extracting object properties. Not only does FlexCap allow conditioning on the box and the length of the answer, it also allows textual prompting of a caption. We can use this to our advantage to extract object properties of interest such as color and material. We show examples of this in Fig. 7. Say we want to extract the color and material of objects, we design two prefixes: 1) "LENGTH-4 The color is ____" and 2) "LENGTH-5 This is made of ____". The pre-fixed length tokens guide the model to produce outputs that are just one word. Note how the model is aware of the object in the bounding box and completes the caption based on the input prefix.
@@
Upon further exploration, we discover prefixes which can extract different kinds of information from an image. We show examples of this in Fig. 8. FlexCap can extract the following attributes: the actions of human ("The person is"), the uses of objects ("This is used for"), OCR ("The sign says"), names of books ("The book is called"), authors of books ("Written by"), and the locations where the photos were taken ("The photo was taken in"). Note how the same model is able to extract different information from the same bounding box: the name of the book title and the author name based on the prefixing (images 4 and 5). In Fig. 9, we highlight how different prefixes can lead to generation of diverse captions. Note how the model correctly identifies the country, possibly using the logo on the train. We find the prefix "Notice" leads to captions highlighting noteworthy aspects in an image which the un-prefixed caption does not. We show more examples of this prefixing in the supplementary material.
Zero-shot Dense Captioning. We can use FlexCap with OWL-ViTv2 [34] to perform dense captioning without further training. We use OWL-ViTv2 to propose bounding boxes in the image. FlexCap then labels each bounding box with a caption. We show examples of this in the project webpage.
5.5 Ablations and Analysis
@@
Model Scaling. Next, we measure the impact of changing the image encoder’s size. We train two models: ViT-B/16 (85M) v/s SO-ViT/14 (428M). Just like the data scaling experiments, both of these have been pre-trained on the WebLI dataset in a contrastive manner. The text decoder is kept the same architecture for both these experiments. We report the results of this experiment in Table 3(c). We find that the larger model results in better performance.
@@
Table 4: Ablations. We vary different aspects of model pre-training, dataset size, and model size and measure its effect on region capitoning task on the Visual Genome dataset. In (a) and (b) the visual backbone is ViT-B/16 and in (b) and (c) all models start from contrastively pre-trained weights.
Compliance. We observe that our model produces captions in a controllable manner with minor deviations. In Figure 10, we show qualitative examples of the FlexCap model producing different length captions for the same box. Note how the model progressively adds more information about the object by incorporating context in the longer sentences (in the jungle), attributes (pink flamingo kite), and alternative nouns (chevy, feline). In the appendix, we have more analysis on length compliance.
6 Related Work
Visual question answering (VQA), a task designed to assess if a computer can answer questions about an image, often requires grounding visual concepts and reasoning. Although initially introduced for supervised evaluation of the task [5], most recently VQA also become one of the most powerful benchmarks for evaluating task and dataset independent visual dialog. Several existing models such as ViperGPT [43], Flamingo [2], BLIP [29, 28], PaLI [8] show convincing zero-shot performance on the VQA benchmarks that rivals the supervised approaches. Unlike most previous zero-shot approaches, which tightly couple vision and language components in a single model, FlexCap generates a high-level human interpretable representation of an image and demonstrates that, through straight-forward application of LLMs, we can achieve comparable performance with state-of-the-art results across VQA benchmarks. Unlike others, ViperGPT [43], also decouples vision and language components and reinterprets visual questions with LLM generated programs, and executes them using existing visual perception tools. Whereas, in our case we use only one powerful vision tool, i.e. FlexCap, to generate all the necessary information and leave the reasoning to an LLM. In that sense, FlexCap is quite complementary to ViperGPT as it can be one of the powerful tools that can improve the controllable visual understanding of the image for ViperGPT.
Length-controlled image captioning has been explored in ZeroCap [45] and LIC [11]. ZeroCap [45] implements length control as a post-processing step by changing the probability of sampling the end-of-sentence token. Hence the model is not naturally trained with word length conditioning in mind and cannot guarantee fine-grained length control at the level of number of words. On the other hand, LIC [11] generates length-controllable captions by conditioning the model with learned tokens that represent different length intervals. However there are considerable differences compared to FlexCap. First, our approach allows for controllability at the level of image regions, while LIC only provides full image captions. This is a significant difference, as it allows us to generate concise or detailed captions for all the objects in the image. Second, our approach has a more precise level of caption-length control. LIC uses a coarse subjective level of control with four or five levels of length (e.g. short, medium, long, and longer), while our approach allows for an exact number of words to be specified.
7 Conclusion
In this work, we introduce FlexCap, a flexible captioning model that can describe localized regions in an image with controllably rich captions. To train FlexCap, we generate a large-scale image-box-caption dataset that is rich in diversity of visual descriptions and their length. We achieve it by utilizing existing web-scale noisy image-text pairs and open-vocabulary object detection models. We show how localized rich descriptions provided by FlexCap can help us connect images and videos to LLMs and achieve strong performance on visual question answering and dense captioning tasks. We also show that our FlexCap model benefits from contrastive pretraining, localized captions dataset size scaling, and model size scaling, and is compliant to the length conditioning. We also demonstrate the effectiveness of FlexCap-enabled localize-then-describe approach over the describe-then-localize approaches.
Acknowledgements
Appendix
Figure 11: Interactive demonstration of FlexCap.
(a) WebLI Dataset
(b) YFCC100M Dataset
Appendix 0.A Interactive Demonstration of Dense Captioning using FlexCap
Through the project webpage https://flex-cap.github.io/flexcap-spatial.html, we demonstrate interactively how well flexible captioning works qualitatively on 40 images randomly selected from the Visual Genome dataset. The interface is shown in Fig. 11. Once the user clicks an image on the webpage, it shows the top-5 bounding boxes by default, but the user can draw any box to get a description for it using FlexCap. Since we cannot pre-compute all potential bounding boxes, we used OWL-ViTv2’s box proposal network to identify 200 boxes and captioned them with FlexCap.
The interactive interface allows a user to draw a bounding box and then the closest pre-computed box is displayed if the IOU between the two boxes is above a certain threshold. We hope this interactive demonstration shows the capabilities of the FlexCap model in detecting and describing different objects and regions in an image.
Appendix 0.B Localized Captions Dataset Details
WebLI. Our dataset creation process on the WebLI dataset resulted in $\sim$32 billion Image-Box-Caption triplets. In Figure 11(a), we show the distribution of caption lengths in the generated dataset using WebLI. We observe that the distribution is not uniform. This is due to the fact that there are more n-grams of length $1$ to sample than length $8$. The average number of unique boxes in an image is $4.19$, and the average number of captions per box is $4.04$.
YFCC100M.Our dataset creation process on the CLIP subset of the YFCC100M dataset resulted in $\sim$0.2 billion Image-Box-Caption triplets. In Figure 11(b), we show the distribution of caption lengths in the generated dataset using YFCC100M. The average number of unique boxes in an image is $6.3$, and the average number of captions per box is $3.3$.
Figure 14: Evaluating open-vocabulary outputs from FlexCap with the help of the CLIP [38] text encoder.
Desired Mean of Std. Dev. of Accuracy Desired Mean of Std. Dev. of Accuracy
Length Pred. Length Pred. Length Length Pred. Length Pred. Length
1 1.02 0.19 0.99 5 5.06 0.26 0.94
2 2.05 0.24 0.96 6 6.01 0.20 0.97
3 3.06 0.29 0.95 7 7.02 0.22 0.95
4 4.04 0.23 0.96 8 8.02 0.23 0.95
Table 5: Compliance metrics. FlexCap produces length-compliant captions for different lengths.
Appendix 0.C Compliance Metrics
In this experiment, we measure how well our model complies to the desired caption length. To do so, we take 1000 images from the MS-COCO dataset and use a random object in the image to produce a description with different lengths. We report the average length of the predicted caption, standard deviation of the length of the predicted caption, and fraction of times the predicted caption has a length equal to the desired length in Table 5.
VOC @ 224 VOC @ 448 COCO @ 224 COCO @ 448
No. of Captions Acc. mAP Acc. mAP Acc. mAP Acc. mAP
1 86.0 84.2 93.2 90.7 71.4 63.2 78.4 72.0
2 86.9 85.2 93.6 91.2 72.0 64.2 79.1 72.9
5 89.3 88.9 95.1 93.3 74.6 69.0 81.3 76.5
10 91.5 93.0 97.2 96.3 77.0 74.9 84.5 82.0
15 92.0 93.7 97.3 96.6 77.8 76.5 85.5 83.8
20 92.3 94.3 97.3 97.1 78.5 77.7 86.0 85.0
Table 6: Object Classification. Accuracy and mAP increase considerably with number of captions generated by the model.
Appendix 0.D Object Classification
Dataset and Evaluation Metrics. In this experiment, we solely evaluate the recognition capabilities of our model in a zero-shot manner. We evaluate how good our model is at recognizing objects at different scales and under occlusion in object detection datasets. To do so, we consider the setting where an image and the ground truth bounding box is provided as input to our model such that it produces a short description of what is contained in the bounding box. For each box, we generate 20 captions each for 4 caption lengths: 1, 2, 3, 4. That is in total, we generate 80 captions. We use 2 object detection datasets: PASCAL-VOC [13] and MS-COCO [31] for this experiment. For each of these datasets, we present 2 metrics: the classification accuracy and the mean Average Precision (mAP) over all ground truth classes.
Mapping Predicted Captions to Classnames. Since our model produces captions for each box, we need to map the predicted descriptive sentences to classnames. To do so, we use an off-the-shelf text encoder (CLIP’s text encoder [38]) to match the generated captions with known object class names. We show the full pipeline in Figure 14. We then take the multiple predicted captions and the ground truth class-names and pass both of these lists through the same text encoder. We then match the text embeddings of the predicted captions with the class names. To do so, we take the mean of the text embeddings of the predict captions. This serves as a query embedding. We then take the cosine similarity of the query embedding and the class name embeddings to produce classification scores. We report the results of this experiment in Table 6. Interestingly, the more descriptions our model generates the more accurate the class predictions become. At a resolution of $224\times 224$, we find mAP improvements of 10.1% (VOC) and 14.5% (COCO) by just generating more caption samples per bounding box. At a resolution of $448\times 448$, we find mAP improvements of 6.4% (VOC) and 13% (COCO). Similar to how using multiple text prompts (a photo of, a drawing of the) with the same object name can lead to a more robust score for different class names (as shown in CLIP [38]), we find that describing the same object with different captions can lead to more robust object classification. Note that these performance gains are training free.
Appendix 0.E Open-Ended Object Detection

### v2

@@ 4.2 Visual Question Answering
FlexCap-LLM. In Figure 1 (right), we show how we use FlexCap with an LLM to solve visual questions. First, we convert an image to a sequence of localized descriptions that describe the image in terms of the objects and regions present in the image. To do so, we need region proposals. We use OWL-ViTv2 [35] to localize important objects and regions in an image. We keep the top 128 bounding boxes by their objectness scores. We then use FlexCap to describe each box in the image in the context of the entire image. In order to produce holistic descriptions, we use multiple prefixes for each region. These prefixes are a combination of length conditioning token and some initial text. We add the boxes and their descriptions to a text preamble as context to the LLM (see Figure 1) that defines the setup where we are using an LLM to answer questions about an image. In all the experiments, we use PALM2-S model [4] as the LLM of choice. We refer to this end-to-end system that takes an image and a question to output the answer as FlexCap-LLM.
Hence we evaluate the effectiveness of FlexCap-LLM on several image VQA benchmarks such as OKVQA [34], VQAv2 [16], GQA [23], and VizWiz [18], and video question answering benchmarks such as MSRVTT [55] and MSVD [54]. Diverse characterictics of these datasets helps gaining better insight on FlexCap’s capabilities. We report the commonly used accuracy metric for each dataset.
Method Acc.(%) $\uparrow$
PalmE-562B [14] 80.0
VLMO [6] 82.8
BEIT-3 [49] 84.2
PaLI-17B [9] 84.3
FewVLM [25] 47.7
Flamingo [2] 56.3
BLIPv2 [29] 65.2
FlexCap-LLM 65.6
(a) VQAv2 results on test-dev set
@@
InstructBLIP [11] 49.5
(b) GQA results on test-dev balanced set
@@
InstructBLIP [11] 34.5
FlexCap-LLM 41.8
(d) VizWiz results on test-dev set
@@
Method MSRVTT MSVD
@@
Table 3: Zero-shot video question answering results reported on MSRVTT-QA and MSVD-QA on the test set. FlexCap-LLM is better than other zero-shot baselines for video VQA benchmarks.
@@
Image Question Answering. First we evaluate FlexCap-LLM on VQAv2 , GQA, OKVQA and VizWiz image VQA benchmarks in a zero-shot setting, meaning that our approach is not trained with the task or the corresponding dataset. The results on these benchmarks are presented in Table 2.
Standard VQA. The VQAv2 dataset is a standard for evaluating the performance of visual question-answering systems. In Table 2(a), we present the results of our evaluation of FlexCap-LLM on this dataset. We find that FlexCap-LLM outperforms other zero-shot baselines, such as BLIP-2. This performance is achieved by providing object and region level information to LLMs without requiring multi-modal fine-tuning.
Compositional VQA. The GQA dataset is for evaluating the performance on complex compositional questions. As FlexCap produces information for multiple visual elements in the scene with their corresponding locations, it is quite well-suited for questions on compositional understanding of the image. On this benchmark, as shown in Table 2(b), FlexCap-LLM outperforms all the recent baselines except InstructBLIP [11].
VQA with External Knowledge. OKVQA dataset is particularly designed for evaluating the ability to answer questions about images that require external knowledge which is not readily available on the image. Hence it requires multiple levels of understanding of the content, and reasoning with that information, which is well-suited for applying FlexCap. In Table 2(c) we show our performance on OKVQA is superior to strong baselines such as Flamingo and ViperGPT which highlights the effectiveness of the mix of generic and specific descriptions generated by FlexCap. Unlike other baselines which use the question, FlexCap generates captions without having access to the question.
VQA with atypical images. We also evaluate on VizWiz, which contains visual questions asked by people who are visually impaired. Unlike web content, in these images the objects and the scene are not always well-centered, hence this dataset contains many out-of-distribution samples compared to typical web-crawled datasets. We report the results of this experiment in Table 2(d). Nevertheless, our approach significantly outperforms Flamingo [2] and InstructBLIP [11] in the zero-shot setting.
Video Question Answering. We also evaluate FlexCap-LLM on zero-shot video question answering datasets MSRVTT-QA and MSVD-QA [54]. The results on these benchmarks are presented in Table 4. For processing the video, we sample 8 frames uniformly from the video. We pass each of these frames through FlexCap to produce captions of objects and regions. We then combine all the object captions from the different frames into one prompt for the LLM. We observe FlexCap-LLM significantly exceeds the performance of the Flamingo model in the zero-shot setting. These results highlight the zero-shot effectiveness of our method, which can solve tasks in the video domain even though both FlexCap and the LLM were not trained for those tasks.
4.3 Dense Captioning
Fine-tuning FlexCap. We fine-tune the pretrained FlexCap model on the Visual Genome train split for 60k steps with a lower learning rate of $61e-$ 6 at a resolution of $448\times 448$.
Captioning GT boxes. Following the evaluation procedure from [26], we evaluate captioning of the ground-truth boxes in Visual Genome. Since this setting removes the localization task, we have a cleaner evaluation of only the region captioning problem. The results of this experiment are provided in Table 4(a) in which we show that FlexCap achieves better performance compared to other approaches evaluated in this setting.
Captioning GRIT boxes. In this experiment, we want to compare against other approaches that perform both localization and captioning. We are measuring how well FlexCap performs when provided with box proposals. Table 4(b) shows FlexCap obtains better performance compared to other approaches evaluated in this setting. Since in our work we do not propose any localization module, we use GRIT’s [52] region proposals as the input boxes for our model. This also allows us to directly compare our captioning capabilities against GRIT. We find that our approach outperforms GRIT at this task even though we test at a lower resolution of $448\times 448$.
@@ 4.4 Open-Ended Object Detection
Open-ended object detection, that is identifying all the objects in the image with rich descriptions, can be achieved in two ways: (a) localize-then-describe or (b) describe-then-localize. With FlexCap, we employ a localize-then-describe approach where boxes are generated with a box proposal mechanism (i.e. OWL-ViT) and then described by FlexCap. We compare these results with a describe-then-localize approach where we first ask a state-of-the-art (SOTA) VLM (i.e. LLAVA [33]) for a comprehensive description of the image with all objects and then localize these descriptions using a SOTA open-vocabulary object detection method (i.e. OWL-ViT [36]). The Recall of all the objects in the image is a major criteria in open-ended object detection. As shown in Figure 6(a), we demonstrate that localize-then-describe approach powered with FlexCap performs significantly better compared to describe-then-localize approach using existing works. In Figure 6 (b-e), we observe that our localize-then-describe approach retrieves more parts of the image, particularly the small and medium size objects, compared to the describe-then-localize approach. This is mainly because most VLMs are trained to describe the salient objects in the image rather than exhaustively list all the objects present in the scene.
@@
4.5 Discussion
@@
Prefixing. Training on a large dataset also enables the extraction of different kinds of information from an image using prefixes. We can use this property to our advantage to extract object properties such as color and material. We show examples of this in Figure 7. Say we want to extract the color and material of objects, we design two prefixes: 1) "LENGTH-4 The color is ____" and 2) "LENGTH-5 This is made of ____". The pre-fixed length tokens guide the model to produce outputs that are just one more word. Note how the model is aware of the input prefix and completes based on the object in the bounding box without confusing it with other surrounding objects.
@@
We show more examples of using the prefix to define in Fig. 8. FlexCap can extract the following attributes with the corresponding prefix in the brackets: the actions of humans ("The person is"), the uses of objects ("This is used for"), OCR ("The sign says"), names of books ("The book is called"), authors of books ("Written by"), and the locations where the photos were taken ("The photo was taken in"). Note the same model can extract different information from the same bounding box: the name of the book title and the author name based on the prefixing (col 4-5).
Limitations. We use OWL-ViT and alt-text to generate the training dataset for the model rather than relying on human annotations. While this approach allows us to scale, there exist biases in the model and alt-text which will reflect in the outputs of FlexCap. Another limitation is that the proposed FlexCapLLM system is not end-to-end trainable but is a composition of a VLM and LLM. This can be alleviated by training a VLM with the localized captioning dataset.
Broader Impact. Localized and controllable image captioning enabled by our model may benefit applications like accessibility tools for visual impairments and intuitive human-computer interaction. However, biases in training data raise concerns around potential misrepresentation or exclusion of certain demographics. Dual-use risks also exist if this model is employed for unethical surveillance.
5 Related Work
Visual question answering (VQA), a task designed to assess if a computer can answer questions about an image, often requires grounding visual concepts and reasoning. Although initially introduced for supervised evaluation of the task [5], most recently, VQA also become one of the most powerful benchmarks for evaluating task and dataset independent visual dialog. Several existing models such as ViperGPT [44], Flamingo [2], BLIP [30, 29], PaLI [10] show convincing zero-shot performance on the VQA benchmarks that rivals the supervised approaches. Unlike most previous zero-shot approaches, which tightly couple vision and language components in a single model, FlexCap generates a high-level human interpretable representation of an image and demonstrates that, through straight-forward application of LLMs, we can achieve comparable performance with state-of-the-art results across VQA benchmarks. Unlike others, ViperGPT [44], also decouples vision and language components and reinterprets visual questions with LLM generated programs, and executes them using existing visual perception tools. Whereas, in our case we use only one powerful vision tool, i.e. FlexCap, to generate all the necessary information and leave the reasoning to an LLM. In that sense, FlexCap is quite complementary to ViperGPT as it can be one of the powerful tools that can improve the controllable visual understanding of the image for ViperGPT.
Length-controlled image captioning has been explored in ZeroCap [46] and LIC [13]. ZeroCap [46] implements length control as a post-processing step by changing the probability of sampling the end-of-sentence token. Hence the model is not naturally trained with word length conditioning in mind and cannot guarantee fine-grained length control at the level of number of words. On the other hand, LIC [13] generates length-controllable captions by conditioning the model with learned tokens that represent different length intervals. However there are considerable differences compared to FlexCap. First, our approach allows for controllability at the level of image regions, while LIC only provides full image captions. This is a significant difference, as it allows us to generate concise or detailed captions for all the objects in the image. Second, our approach has a more precise level of caption-length control. LIC uses a coarse subjective level of control with four or five levels of length (e.g. short, medium, long, and longer), while our approach allows for an exact number of words to be specified. [48] also propose an approach to produce variable length descriptions for objects localized interactively. They use a pre-trained image captioner to produce descriptions of objects and ChatGPT in post-hoc to output length-conditioned captions. While in our FlexCap model, the length tokens modulate the output produced by the captioner.
6 Conclusion
In this work, we introduce FlexCap, a flexible captioning model that can describe localized regions in an image with controllably rich captions. To train FlexCap, we generate a large-scale image-box-caption dataset that is rich in diversity of visual descriptions and their length. We achieve this by utilizing existing web-scale noisy image-text pairs and open-vocabulary object detection models. We show how localized rich descriptions provided by FlexCap can help us connect images and videos to LLMs and achieve strong performance on visual question answering and dense captioning tasks. We also show that our FlexCap model benefits from contrastive pretraining, localized captions dataset size scaling, model size scaling, and is compliant to length conditioning. We also demonstrate the effectiveness of FlexCap-enabled localize-then-describe approach over the describe-then-localize approaches.
Acknowledgements
Appendix
Appendix A Localized Captions Dataset Details
WebLI. Our dataset creation process on the WebLI dataset results in $\sim$32 billion Image-Box-Caption triplets. In Figure 9(a), we show the distribution of caption lengths in the generated dataset using WebLI. We observe that the distribution is not uniform. This is due to the fact that there are more n-grams of length $1$ to sample than length $8$. The average number of unique boxes in an image is $4.19$, and the average number of captions per box is $4.04$.
YFCC100M.Our dataset creation process on the CLIP subset of the YFCC100M dataset results in $\sim$0.2 billion Image-Box-Caption triplets. In Figure 9(b), we show the distribution of caption lengths in the generated dataset using YFCC100M. The average number of unique boxes in an image is $6.3$, and the average number of captions per box is $3.3$.
(a) WebLI Dataset
(b) YFCC100M Dataset
Appendix B Ablations and Analysis
@@
Model Scaling. Next, we measure the impact of changing the model size. We train two models: ViT-B/16 (85M) v/s SO-ViT/14 (428M). Just like the data scaling experiments, both of these have been pre-trained on the WebLI dataset in a contrastive manner. We report the results of this experiment in Table 5(c). We find that the larger model results in better performance.
@@
Table 5: Ablations. We vary different aspects of model pre-training, dataset size, and model size and measure its effect on region captioning task on the Visual Genome dataset. In (a) and (b) the visual backbone is ViT-B/16 and in (b) and (c) all models start from contrastively pre-trained weights.
Appendix C Open-Ended Object Detection